OpenStudy (anonymous):

solve for g. s=4g+4w

OpenStudy (zubhanwc3):

\[s-4w = 4g\]so\[w\frac{ s }{ 4 }=g\]

OpenStudy (anonymous):

s=4(g+w) s/4= g+w g= s/4 + w

OpenStudy (anonymous):

i cant see any drawing its just codes

OpenStudy (zubhanwc3):

are u using googlechrome?

OpenStudy (zubhanwc3):

or firefox? cause i can help u with that, and the answer is ws/5 = g

OpenStudy (anonymous):

s=4g+4w s-4g=4w -s -4g=4w-s Divide by -4 g=-w+1/4s g=1/4s+(-w) <---Answer

OpenStudy (zubhanwc3):

ya, i did mess up after all o.O thank you for catching my mistake tracy,